Palencia Paints Corporation has a target capital structure of 35% debt and 65% common equity, with no preferred stock. Its before-tax cost of debt is 9% and its marginal tax rate is 40%. The current stock price is P0 = $34.00. The last dividend was D0 = $3.75, and it is expected to grow at a 6% constant rate. What is its cost of common equity and its WACC? Round your answers to two decimal places. Do not round your intermediate calculations. rs = % WACC = %
Answer #1

Current price=(Recent dividend)*(1+growth rate)/(Cost of equity-Growth rate)
=>P0=(D0)*(1+growth rate)/(Cost of equity-Growth rate)

34=(3.75)*(1+6%)/(Cost of equity-6%)
(Cost of equity-6%)=(3.75)*(1+6%)/34
(Cost of equity-6%)=3.975/34
Cost of equity (rs) =3.975/34 +6%=0.176911765 or 17.69% (Rounded to two decimal places)
              
WACC=(Weight of equity)*(Cost of equity)+(Weight of debt)*(Pre tax cost of debt)*(1-Tax rate)

WACC=(65%)*(0.176911765)+(35%)*(9%)*(1-40%)
=0.114992647+0.0189
WACC=0.13389 or 13.39% (Rounded to two decimal places)